2. The molar solubility of PbBr is 2.17 x 10 M at a certain temperature. Calculate Kp for PbBr. (7)Please show step-by-step work. Thank you!
Answer #1

The PbBrz is soluble in hot water.Its solubility equilibrium is given by-

PbBr_{2}_{(s)}\rightleftharpoons Pb^{2+}_{(aq)}+2Br^{-}_{(aq)}

Now the Solubility product constant of the above solubility equilibrium is given by-

K_{sp}=[Pb^{2+}].[Br^{-}]^{2}---1

If the Molar Solubility of the salt is 'S" (say),then from equation 1,we get,

K_{sp}=S.(2S)^{2}=S.4S^{2}=4S^{3}

Given that, S = 2.17 × 10-3M ,we get-

K_{sp}=4.(2.17\times 10^{-3})^{3}=4.09\times 10^{-8}.